LO 2.3Markson and Sons leases a copy machine with terms that include a fixed fee each month of $500 plus a charge for each copy made. The company uses the high-low method to analyze costs. If Markson paid $360 for 5.000 copies and $280 for 3,000 copies how much would Markson pay if it made 7.500 copies ?
Answer #1

Correct answer-----------It will pay $460

Working

Cost No. of activities
A High Level $                360.00 5000
B Low Level $                280.00 3000
C=A-B Difference $                   80.00 $              2,000.00
A Cost difference $                   80.00
B No. of activities difference 2000
C=A/B Variable cost per unit $                     0.04
High Level Low Level
A No. of activities 5000 3000
B Variable cost per unit $                0.0400 $                  0.0400
C=AxB Total Variable cost $                200.00 $                  120.00
D Total cost $                360.00 $                  280.00
E=D-C Total fixed cost $                160.00 $                  160.00

.

No of Copies 7500
Variable cost $           300.00
Fixed cost $           160.00
Total cost $           460.00